The time taken to complete a web-based questionnaire is normally distributed, with an average time (μ) of 9 minutes and a standard deviation (σ) of 1.55 minutes. What is the probability that a randomly selected person will take:
1
Expert's answer
2021-10-25T16:56:55-0400

Let "X=" the time taken to complete a web-based questionnaire: "X\sim N(\mu, \sigma^2)."

Given "\mu=9\ min, \sigma=1.55\ min."


"P(8<X<8.5)=P(X<8.5)-P(X\leq 8)"

"=P(Z<\dfrac{8.5-9}{1.55})-P(Z\leq \dfrac{8-9}{1.55})"

"\approx P(Z<-0.32258)-P(Z\leq -0.64516)"

"\approx0.3735064-0.2594113=0.114095"

The probability that a randomly selected person will take between 8 and 8.5 minutes to complete the questionnaire is "0.114095."




"P(8.75<X<9.75)=P(X<9.75)-P(X\leq 8.75)"

"=P(Z<\dfrac{9.75-9}{1.55})-P(Z\leq \dfrac{8.75-9}{1.55})"

"\approx P(Z<0.48387)-P(Z\leq -0.16129)"

"\approx0.6857613-0.4359325=0.249829"

The probability that a randomly selected person will take between 8.75 and 9.75 minutes to complete the questionnaire is "0.249829."